OpenStudy (anonymous):

2x² + 5x = 12? Do i just move equal sign down past 12 and then make the whole equation equal zero?

hartnn (hartnn):

yes, you can do that by subtracting 12 from both sides.

OpenStudy (anonymous):

so 2x² + 5x +12= 0 Then subtracting 12 2x^2 +5x= -12. I'm sorry I am way off

hartnn (hartnn):

like this : 2x² + 5x = 12 subtracting 12 from both sides 2x² + 5x-12 = 12-12 2x² + 5x -12=0 got this ?

OpenStudy (anonymous):

yes ok

hartnn (hartnn):

now do you know how to solve quadratic equations ?

hartnn (hartnn):

by factoring ?

OpenStudy (anonymous):

Can you walk me through it?

hartnn (hartnn):

sure, in general for \(ax^2+bx+c=0\) to factor, we search 2 numbers whose product is a*c and sum is 'b' so, for your Q, 2x² + 5x -12=0 can you think of 2 numbers with product -24 and sum = 5 ?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

and um, no I cant think of the other. Im clueless

hartnn (hartnn):

you got one number as 8 ! thats correct, now for sum to be 5, the other number must be 5-8 =-3, right ?

hartnn (hartnn):

check, product = 8*(-3) =-24 = correct.

OpenStudy (anonymous):

ah damn, i was going to say that! lol

hartnn (hartnn):

good, so now you can split the middle term as 5x = 8x-3x so, 2x^2+8x-3x-24=0 got this step ?

OpenStudy (anonymous):

oh wow, yes okay I am writing this all down, but yes

OpenStudy (anonymous):

This is all great thanks so much!

hartnn (hartnn):

welcome ^_^ but we are not done yet... :) and sorry, it was 2x^2+8x-3x-12=0 (not -24) next thing is to factor out from 1st 2 terms : 2x^2+8x , factor out 2x, what u get ?

hartnn (hartnn):

you there ? having any trouble? or doubts?

OpenStudy (anonymous):

Sorry! My computer died on me last night and I had to fully charge it over night!

OpenStudy (anonymous):

2x(x + 4) - 3(x + 4) = (2x - 3)(x + 4) ??

hartnn (hartnn):

thats correct :)

hartnn (hartnn):

now do you have to solve for x ? or just factorize ?